### v2

# Quantifying urban socio-economic segregation through co-residence network reconstruction

Abstract
Urban segregation poses a critical challenge in cities, exacerbating inequalities, social tensions, fears, and polarization. It emerges from a complex interplay of socio-economic disparities and residential preferences, disproportionately impacting migrant communities. In this paper, using a comprehensive administrative data from Vienna, where nearly 40% of the population consists of international migrants, we analyse co-residence preferences between migrants and locals at the neighbourhood level. Our findings reveal two major clusters in Vienna shaped by wealth disparities, district diversity, and nationality-based homophily. These insights shed light on the underlying mechanisms of urban segregation and designing policies for better integration.
Introduction
Cities offer unparalleled access to essential services, including healthcare, education, transportation, cultural amenities, water and sanitation services (Glaeser and Maré,, 2001; Leon,, 2008; Winters,, 2011). Cities are the hubs where people have increased social interactions, fostering productivity and attracting talent (Keuschnigg et al.,, 2019). Some of those positive urban features are further accentuated in large cities. For example, large cities tend to pull more people, primarily young (Prieto-Curiel et al.,, 2022) and have a much bigger share of international migrants (Prieto-Curiel et al.,, 2018). Although large cities are an ideal space for social mixing and cultural integration, they can also form highly unequal and segregated communities, where groups rarely interact with others (Sousa and Nicosia,, 2022). In some cities, for example, wealthy neighbourhoods tend to segregate themselves from the urban population and build themselves private cities (Pieterse et al.,, 2018). Segregated cities often result from the residential selection process, where even small preferences for some neighbour characteristics highly alter the distribution of population groups (Clark and Fossett,, 2008; Bolt et al.,, 2008; Lee,, 2016).
The emergent phenomenon of segregation greatly impacts the living spaces and design planning of cities (Bruch and Swait,, 2019). In principle, most cities are not designed to be segregated, but individual residential selections, governmental programs (Trounstine,, 2018), mobility patterns (Candipan et al.,, 2021), and preferences to live near similar others may result in segregated cities (Clark and Fossett,, 2008). On the positive side, segregation has been studied as a mechanism that communities use to create and preserve social capital (Bourdieu,, 1986). Fueled by trust, mutual understanding, shared values, and behaviours, the social capital coming from segregation aims for cultural preservation for immigrant and local communities and to maintain socio-economic positions (Portes and Zhou,, 1993; Light and Thomas,, 2019; Pieterse et al.,, 2018). However, high levels of economic, social, and geographical segregation accentuate social, health, and education inequalities (DeVerteuil,, 2009). Underprivileged communities are at risk of isolation, marginalization, and victimization (O’Flaherty and Sethi,, 2007; Desmond and Shollenberger,, 2015) and intergroup conflict can emerge from hostility and cultural misunderstandings (Pettigrew and Tropp,, 2006; Helbing et al.,, 2015). The interplay between those different dimensions of segregation creates the need for global cities to avoid becoming disintegrated ghettos and tackle the effect of segregation on increasing polarisation, political extremism, and perception biases towards minorities (Düring and Wolfram,, 2015; Motyl et al.,, 2014; Lee et al.,, 2019).
Vienna is the capital, largest city, and one of nine federal states of Austria, with about two million inhabitants spread among its 23 districts (Statistik Austria,, 2024). Its central position makes Vienna a key migration hub in Europe and a unique case for studying segregation patterns, particularly due to its historical context and its distinction as a leading city for quality of life (Economist-Intelligence,, 2024), achieved through policies promoting affordable housing and social integration. In 2023, nearly 40% of Austria’s migrants—defined as individuals with foreign nationalities—resided in Vienna, highlighting the city’s importance as a destination for international migrants (Statistik Austria,, 2024). This is especially true for refugees, who often choose to resettle in Vienna due to the city’s established support systems and services for asylum seekers (Ali et al.,, 2024). Vienna also stands out for its extensive history of social housing, with one of the largest stocks of community-owned buildings in Europe (Kohlbacher and Reeger,, 2016; Social Housing Vienna,, 2024). This combination of a high migrant concentration and a strong tradition of public housing creates a unique environment in which the dynamics of integration and segregation can be explored in depth. However, despite the city’s high level of diversity, the extent of segregation within Vienna remains unclear. A comprehensive understanding of how citizens from various backgrounds integrate is still lacking. While some studies suggest spatially segregated patterns of migration (Kohlbacher and Reeger,, 2016, 2020), the whole dynamics of residential integration have yet to be fully understood.
To investigate these dynamics, we use granular administrative data related to all migrants in Austria. The data includes all migrants in Austria who register their residence through the mandatory registration form called the “Meldezettel”. By applying data mining and statistical analysis, we develop a metric to quantify co-residence preferences among populations from diverse nationalities. This measure allows us to construct a global network linking countries based on significant shared residence priorities and identify two prominent clusters with similar co-living patterns.
Results reveal that in Vienna, disparities in neighbourhood wealth and the intrinsic diversity of districts play a critical role in shaping co-residence ties. Furthermore, we observe that individuals’ preferences to live among others from similar backgrounds (homophily) significantly influence the formation of clusters. These factors act as both attractors and deterrents, shaping residential decisions and contributing to the broader patterns of segregation in Vienna.
Results
Employing a comprehensive dataset of residence registrations of migrants in Vienna, this study investigates how the co-residence of people from diverse nationalities in the city can contribute to the emergence of urban socio-economic segregation. We employ data mining and statistical analysis techniques to develop a metric of co-residence preferences among migrants and local populations and use this measure to construct a world map of residence preferences.
Data description.
Our primary data source is provided by the Austrian government, offering accurate residential data for all migrants residing in Austria since November 2022. Each migrant in the dataset is characterised by their foreign citizenship, which is obtained from the mandatory residence registration form (“Meldezettel”). We used a snapshot of the residence locations of the entire migrant population in the districts of Vienna captured on 22nd September 2023 (Fig. 1a). To ensure data privacy, residence information is aggregated and includes only nationalities represented by more than ten individuals per district. We focused on the 19 most populous migrant nationalities in Vienna, categorising all other nationalities under the label “Others” (Fig. 1b). For further details on the data, see Methods and Supplementary Note LABEL:sec:residencedata. In addition to residential data, complementary datasets from publicly accessible sources are collected to enhance the comprehensiveness of our studies (see Methods and Supplementary Note LABEL:sec:otherdatasets).
Figure 1: Melting Pot in Vienna. a Population of migrants in the districts of Vienna on 22nd September 2023 (in %). Districts 15 and 13 have the highest and lowest percentage of migrants, respectively (see Supplementary Table LABEL:tab:tabs1 for a detailed list of values). b Population distributions of the top 19 most populous migrant nationalities in Vienna represented by at least ten individuals per district, the label “Others” is used for the rest. The colour bar indicates the fraction of people living in Vienna on a log scale (see Supplementary Table LABEL:tab:tabs2 for a detailed list of values). Countries are sorted from highest to lowest population fraction in Vienna.
Mapping co-residence preferences
To explore the boundaries of socio-spatial segregation, we analyse the rates at which individuals from different nationalities reside in the same districts. First, we map all the co-living links between countries based on data on individuals who live in the same districts. Next, using statistical tests, we extract the significance level of each co-residence link, allowing us to remove random associations due to district and population sizes. Finally, we use a network community detection method based on random walks to uncover groups of countries from pairwise connections. This approach provides an interpretable map of the large-scale structure of socio-spatial segregation (Karimi et al.,, 2015).
Extracting co-living links.
In essence, our goal is to establish connections between countries based on whether their citizens reside in the same districts. We first aggregate the total number of individuals from each nationality residing in each district of Vienna (Fig. 1b). When citizens from different countries share a district, we interpret this as those countries having a common interest in co-living within that district. To quantify this, we calculate the probability of a random encounter in the district involving individuals from two specific countries, assuming the district’s population is well-mixed. For a given district $d$, if the population of country $i$ in that district is $\kappa_{i}^{d}$, the metric for district-level co-residence between countries $i$ and $j$ is given by:
$w_{ij}^{d}=\frac{1}{(N^{d})^{2}}\kappa_{i}^{d}\kappa_{j}^{d}\propto\kappa_{i}^{d}\kappa_{j}^{d},$ (1)
where $N^{d}$ is the total population of the district. In practice, we define the co-residence metric as $w_{ij}^{d}=\kappa_{i}^{d}\kappa_{j}^{d}$ and subsequently normalise it using a robust method that accounts for district size. Thus, we approximate the residential preferences of a country by analysing the collective settlement patterns of its citizens residing in the city. The outcome is a large, fully connected network where nodes are countries and links represent co-living between countries.
Extracting significant co-living links.
Using raw co-living counts introduces bias due to the varying size of populations per country and district. This leads to connections driven by large inhabitant communities rather than genuinely shared interests in co-living. To account for this, we employ a statistical validation method that filters out connections likely resulting from size effects or random placement. As illustrated in Fig. 2a, we control for these random effects by simulating random relocations of residents that keep the total populations of countries and districts fixed. We then use the randomised resident counts as a baseline. Suppose the empirical data show that individuals from two countries co-reside in the same district more frequently than expected under the randomised baseline. In that case, it indicates a positive association between the two countries, suggesting a preference or affinity for shared living spaces. Conversely, if the co-residence rate is lower than expected, it points to a negative association, implying a tendency to avoid co-living in the same districts. Formally, this method uses a multinomial distribution to provide analytical expressions for the average and standard deviation of co-living rates, thereby identifying significant links in a bipartite system with countries on one side and districts on the other (Supplementary Fig. LABEL:fig:figs5) (Karimi et al.,, 2015). Other methods exist to evaluate significant correlations between entities in bipartite systems (Zweig and Kaufmann,, 2011; Tumminello et al.,, 2011; Serrano et al.,, 2009). Nonetheless, our model offers the advantage of easily accounting for size variations within districts and across country populations.
To compare empirical and expected link values, we obtain the mean $\mu_{ij}^{d}$ and standard deviation $\sigma_{ij}^{d}$ of the co-residence metrics $w_{ij}^{d}=\kappa_{i}^{d}\kappa_{j}^{d}$ under the randomised baseline for each specific district $d$ and compute z-scores:
$z_{ij}^{d}=\frac{w_{ij}^{d}-\mu_{ij}^{d}}{\sigma_{ij}^{d}}.$ (2)
The z-scores help compare weights, as they account for the significant variations present in the district’s residence patterns. We then aggregate the z-scores over all districts to obtain the cumulative z-score for a pair of countries $i$ and $j$ (see Fig. 2a). Among the 21 groups (19 most populous migrant nationalities, Others, and Austrians), we identified 210 significant links, forming a fully connected, undirected, and weighted (z-scores) network based on districts co-residence. We present the network in Fig. 2b, with 80 positive links shown in red and 130 negative links in blue (see Methods and GitHub repository (Parera,, 2024) for a detailed list of z-score values).
Positive links indicate that, on average, the number of co-living citizens from a pair of countries across all districts exceeds the expected values. This suggests a form of closeness between these countries, reflecting a preference to reside in the same districts. In contrast, negative links suggest an avoidance pattern, where residents from a pair of countries are less likely to live in the same districts than would be expected by random chance. This raises the question of what underlying factors drive these patterns of association. To address this question, we begin by identifying the co-living clusters, filtering out the negative links, and considering exclusively the 80 positive connections.
Clustering countries.
To determine groups of countries with similar residence preferences, we use a community-detection method known as the map equation (Rosvall and Bergstrom,, 2008; Rosvall et al.,, 2009), and its associated search algorithm, Infomap (Edler et al.,, 2023) (see Methods section). Infomap uses the weights of links (z-scores) to determine the clusters of strongly connected countries (see Fig. 2c). To gain better insights into the shape of these two clusters, we examine the country network structure within each cluster (see Fig. 2d and GitHub repository (Parera,, 2024) for a detailed Infomap network). Notably, countries with larger link strengths serve as hubs, connecting to many other countries. These robust connections act as key channels for information exchange, underscoring the significant role these countries play in the dissemination of multicultural information within each cluster. People with the same nationalities may connect based on their shared background, and people living in the same districts may interact due to spatial proximity. Various post-stratification techniques, including implementing the Bonferroni correction, were employed to verify the representativeness and robustness of the findings (see Supplementary Note LABEL:sec:sanity).
Determinants of co-residence clusters
Cultural and national homophily.
The key question is what factors determine the segregation of countries into two clusters. The world map of residence preferences (Fig. 2c) suggests that cultural and geographic proximity could be a key factor in the division of countries (Crang,, 2013). Furthermore, the world network of residence preferences (Fig. 2d) reveals that connections are typically stronger between geographically proximate countries, further supporting this hypothesis. Countries within the same cluster tend to have stronger cultural ties, sharing similar languages, traditions, and social norms. This cultural homophily drives the preference for migrants from these countries to live in close proximity, reinforcing the observed segregation.
The five most populous nationalities in the larger cluster are Austria, Ukraine, Germany, Russia, and Hungary, while the most populous nationalities in the smaller cluster are Serbia, Turkey, Syria, Romania, and Poland. Most countries belonging to the big or small clusters are from culturally similar regions (Crang,, 2013). There are a few notable exceptions. For instance, within the “minority cluster”, Poland and Afghanistan stand out due to their relatively distant geographic locations compared to the other countries in the group. This results in these nations being among the least connected within their cluster, suggesting that other factors may account for their inclusion. Similarly, while China and Iran are part of the “majority cluster”, their cultural and geographic distance from the other member countries leads to weaker connections within the group. This suggests that while cultural homophily is a strong driver of cluster formation, other factors, such as economic reasons, may influence the clustering patterns.
Beyond cultural homophily at the macro-level, national homophily, the tendency for individuals to associate with others from the same nationality, may also significantly contribute to the observed segregation patterns within Vienna. In Methods, we characterise the homophily level for each nationality using the Dissimilarity index, the proportion of members of a group who would need to relocate so that the group fractions at the district-level match the city-wide distribution. Turkey, Germany, and Italy have the highest level of homophily, while Hungary and Slovakia have the weakest level (see Methods section for details).
When a particular nationality exhibits a strong degree of homophily (with the majority of its members residing in the same districts), it can act as an anchor, attracting migrants from other countries with similar cultural backgrounds, forming diasporas (Prieto-Curiel et al., 2024b, ). This concentration not only strengthens the presence of the initial nationality within the area but also encourages nationalities with similar cultural backgrounds to settle nearby, creating pockets of cultural cohesion across the city. Therefore, this micro-level homophily, which captures preferences for co-living within one’s national group, can compound with the clustering effects observed across larger cultural groups. Next, we examine in greater detail how neighbourhood wealth disparities and diversity characteristics also contribute to this clustering pattern, acting as additional forces in shaping co-living preferences.
Figure 2: Clustering countries sharing similar residence interests. a Diagram of the method used to build the network of co-residence. b z-score network. Edges with $\lvert z_{ij}\rvert\leq 20$ are not shown for better readability. The width and colour transparency of the links is proportional to the z-score values, while the size of the nodes reflects the total population of each country, except for Austria. The network has been produced using Noverlap layout of Gephi software (Bastian et al.,, 2009). Two prominent clusters are marked with a dashed line. c World map of residence preferences. Countries that belong to the same cluster have the same colour. Countries coloured with grey lines are either subsumed within Others or for which we lack data on immigrants residing in Vienna. The larger cluster, comprising the majority of the population residing in Vienna, is shown in purple and referred to as the “majority cluster”. The smaller cluster, with fewer residents, is depicted in yellow and termed the “minority cluster”. d World network of residence preferences. The size of the nodes represents the total z-score of the clusters and countries. The links represent the connections between nodes obtained from the cluster analysis with Infomap (Edler et al.,, 2023); the thicker the line, the stronger the connection.
Neighbourhood wealth.
Neighbourhood wealth plays a pivotal role in shaping residential patterns, influencing both where individuals settle and the socio-economic composition of urban areas. By examining the wealth disparities across districts, we can reveal how economic factors drive residential choices, ultimately contributing to the spatial distribution of country populations. To do that, we incorporate an additional publicly available dataset to approximate the wealth of the districts (see the Methods and Supplementary Note LABEL:sec:incomerepresentativeness).
We approximate the wealth distribution at the district level by the average net yearly income of the inhabitants of the district (see the Supplementary Note LABEL:sec:incomerepresentativeness). Then, for each country, we compute the Pearson correlation coefficient between the proportion of the country’s population in a district and the district’s average net income (see Methods section for details). A high positive correlation indicates that the country’s residents are concentrated in wealthier districts, suggesting greater affluence. Conversely, a strong negative correlation implies that its residents tend to live in lower-income districts. This analysis quantifies the tendency of individuals from a specific country to reside in districts where the perceived socio-economic status of other residents or themselves is marked by a higher or lower average wealth of inhabitants. We then rank the countries from highest to lowest Income-Population fraction correlation (Fig. 3a). For instance, Austria exhibits a strong positive correlation, indicating that its residents tend to prefer districts with higher average wealth, reflecting an alignment with higher socio-economic status. In contrast, Bosnia and Herzegovina shows a negative correlation, suggesting that its residents are more likely to live in lower-income districts, potentially influenced by economic constraints or community ties.
There is a clear economic division in Vienna between the majority and the minority clusters. The two identified groups with similar co-living interests are segregated in terms of neighbourhood wealth in Vienna (see Fig. 3a). The analysis indicates that the population from countries belonging to the majority cluster tends to live in districts where residents are wealthier compared to those in the minority cluster, with the exceptions of Slovakia, China, Hungary, and Others, which also exhibit negative correlation values. This disparity in economic status can contribute significantly to the observed segregation of nationalities within the city, where people from the majority cluster tend to live in more affluent districts and vice versa. Supplementary Note LABEL:sec:wealth explores alternative perspectives on neighbourhood wealth disparities, presenting consistent outcomes and complementary analyses based on a dataset of district-level rental prices in Vienna.
Figure 3: Socio-economic and diversity factors underlying nationality clusters. a Correlation between district wealth and population fraction for each nationality. b Correlation between diversity and population fraction for each nationality. Nationalities are ranked by their correlation values, from highest to lowest for wealth-population fraction correlation in a, and from lowest to highest for diversity-population fraction correlation in b. The nationality clusters are colour-coded consistently with Fig. 2c.
Neighbourhood diversity.
Neighbourhood diversity is also strongly correlated with the two clusters. To estimate urban diversity, we introduce the Simpson index, which can be calculated for each district (local) and the whole city. It expresses the probability that two randomly selected individuals from the district (or the city) have different nationalities (see the Methods section for details). After calculating the diversity at the district level, we compute the Pearson correlation coefficient for each country, this time examining the relationship between the country’s population fraction relative to the whole population in a district and the diversity of that district. In this context, the analysis measures the extent to which the population of a specific country experiences diversity within the urban environment. Figure 3b shows that the two identified clusters are distinctly separated with respect to diversity within the districts of Vienna. This observation suggests that populations from countries within the majority cluster tend to reside in districts characterised by lower diversity, resulting in a diminished perception of residing among foreign individuals compared to those in the minority cluster. Countries with correlations close to zero indicate no clear association between diversity and residential patterns, meaning their populations are equally likely to reside in districts with either higher or lower diversity levels.
Diversity within a district may be a consequence of broader socio-economic and cultural factors, such as income, educational background, migrant flows, and geographical proximity. Wealthier and more cosmopolitan districts may naturally attract a lower mix of nationalities due to economic constraints limiting where migrants can afford to live, decreasing the diversity in these areas, while less affluent districts may exhibit higher diversity, resulting in a more heterogeneous population. In this context, the grouping of nationalities into two distinct groups could be due to income inequalities and cultural differences, which are reflected in patterns of residential segregation and district diversity. These associations may introduce biases that further enhance the clustering effect, creating a feedback loop: as certain nationalities settle in particular districts, their presence may influence the perceived diversity of the area, either attracting more migrants from similar backgrounds and reinforcing homogeneity or encouraging further diversity and integration.
Over time, these processes can either sustain or alter the existing patterns of segregation, with district diversity both shaping and being shaped by the clustering tendencies of migrant populations. In this way, biases and self-segregation actively interact with district characteristics, shaping key associations that guide settlement patterns. Observing those patterns requires an investigation over a long period since culture evolves at a faster pace compared to individual movements.
Discussion
Our analysis reveals a pattern of segregation within the Vienna population, characterised by two distinct groups of nationalities. Citizens from the same cluster tend to reside in the same districts, while those from different clusters are less likely to cohabit in shared districts. This clustering can be explained by the interplay between neighbourhood wealth and diversity, with wealth disparities emerging as a key factor. Citizens from wealthier and more educated countries predominantly settle in more affluent districts, reinforcing these patterns of segregation. At the same time, the diversity of the districts reinforces these patterns by encouraging residents to settle in areas with less or more diversity. These factors act as both pull and push mechanisms that influence the residential decisions of migrants in Vienna.
The relationship between district diversity and the clustering of country populations in Vienna is a complex and interdependent process, where diversity could act both as a cause and a consequence of segregation, generating a multifaceted feedback loop. On the one hand, diversity can influence the residential choices of different nationalities, with some inhabitants being attracted to more diverse areas due to their openness to multicultural environments or social adaptability. In such cases, citizens may gravitate toward districts where multiple ethnicities coexist, viewing this diversity as an opportunity for social integration and cultural exchange. Conversely, other nationalities may prefer less diverse neighbourhoods, seeking the comfort and familiarity of living among people from similar cultural, linguistic, or social backgrounds. This mechanism of self-segregation, or the tendency of individuals to self-select into places based on nationality or culture, drives the formation of more homogeneous clusters, reinforcing patterns of low diversity in certain districts. On the other hand, diversity within a district may also be a consequence of broader socio-economic and cultural factors, such as income, educational background, migrant flows, and geographical proximity.
In essence, the observed segregation in Vienna is not solely a product of cultural or economic factors in isolation but rather a result of a dynamic and ongoing interaction between diversity, individual preferences, and systemic influences, all of which contribute to the intricate spatial distribution of migrant communities across the city. This highlights that segregation is not attributable to a singular factor but is instead a multifactorial phenomenon shaped by a multitude of influences. This multifaceted character underscores the inherent complexity behind the residential clustering of citizens in urban areas.
Further work could analyse the temporal evolution of migrant segregation, understanding the shape and scale of the phenomenon of internal migration, particularly toward urbanisation and mobility flows between human settlements with a stochastic Agent-Based model (Zuccotti et al.,, 2023) taking into account the new arrivals depending on the Diaspora model (Prieto-Curiel et al., 2024a, ). Another direction would be to analyse other factors such as urban green areas, municipal housing, Urban Vulnerability index, and the impact of infrastructure on encouraging diversity. Understanding the underlying causes of residential segregation is crucial for developing policies that foster social cohesion and integration. Addressing both economic and diversity factors can help design interventions that promote more social and inclusive neighbourhoods.
Methods
Residence registration data of migrants in Vienna.
The data used here is provided by the Federal Ministry of Interior of Austria, Bundesministerium für Inneres “BMI”. The data includes all individuals in Austria with a foreign nationality (migrants) who register their residence through the mandatory registration form called the “Meldezettel”. The form is updated every time a person changes their primary residence. The dataset thus encompasses all migrants residing in Vienna up until September 2023. The data excludes short-term visitors, such as tourists who are not required to register, and does not account for undocumented migrants (see Supplementary Note LABEL:sec:residencedata).
Residence district locations’ extraction.
To protect individual privacy, population counts are aggregated at the district level, with statistics provided only for nationalities represented by at least ten individuals within a district. By setting this threshold, we mitigate the risk of identifying specific individuals within each district, which is particularly important in small population groups or areas with fewer residents of a given nationality. Aggregating data at this level helps maintain confidentiality by concealing individual characteristics within broader statistical averages. This approach ensures that privacy standards are upheld while still allowing for meaningful insights into district-level patterns and trends in migrant residency. This method identifies the top 19 most populous migrant nationalities residing in Vienna while consolidating the remaining nationalities under the category of “Others”. By focusing on these top 19 countries, we capture the most significant demographic groups contributing to Vienna’s migrant population while ensuring data manageability and privacy compliance. The “Others” category aggregates the less-represented nationalities, allowing us to maintain a comprehensive coverage of all groups in the analysis without risking the identification of individuals from smaller communities.
Complementary datasets.
To support our analyses of the residential patterns of inhabitants in Vienna, we used several additional datasets. The geographic boundaries for the Vienna districts were obtained from the Vienna GIS database (City of Vienna, 2023c, ), while the global GIS data provided country boundaries to facilitate spatial visualisation and cross-national comparisons (Natural Earth,, 2023). Population counts by district were sourced from census data filtering for the year 2023, allowing normalisation of migrant fraction within districts and the extraction of the Austrian national population (City of Vienna, 2023b, ). The average net income data at the district level were filtered for the year 2020 and served as a proxy for economic status, helping to assess the socio-economic factors that influence residential clustering (City of Vienna, 2023a, ). Additional datasets are detailed in the Supplementary Note LABEL:sec:otherdatasets.
Mapping co-residence network.
In this section, we outline the formalisation of our model for measuring the co-residence network between nationalities in the Vienna districts. This network model allows us to identify significant clustering patterns within Vienna’s spatial structure. The connections, or links, between nationalities are based on the degree to which two nationalities co-reside in the same districts (Karimi et al.,, 2015). To compute co-residence links between countries, self-loops, representing co-residence within the same nationality, are excluded from the analysis. This ensures that the resulting network focuses solely on interactions and shared residential patterns between different nationalities, allowing for a clearer investigation of cross-national co-residence dynamics.
We define the empirical link weight $w_{ij}^{d}$ between two nationalities $i$ and $j$ within a specific district $d$ as follows: if nationality $i$ has $\kappa_{i}^{d}$ residents in the district $d$ and nationality $j$ has $\kappa_{j}^{d}$ residents in the same district, then the empirical link weight $w_{ij}^{d}$ is given by:
$w_{ij}^{d}=\kappa_{i}^{d}\kappa_{j}^{d}.$ (3)
This link weight captures the strength of co-residence between two nationalities within a district, representing the potential for interaction or overlap based on residential distribution.
Given that observed link weights may vary widely due to factors such as population size and district-level density, it is necessary to determine the statistical significance of these links. To do so, we compare the empirically observed link weights to a null model. In the null model, we assume that each resident comes from a country randomly picked from the total population of Vienna. The random assignments are performed by drawing the countries’ population within each district from a multinomial distribution. This random distribution accounts for the baseline expectation that two nationalities might co-reside by chance alone, based on their overall presence in the city. From the null model, the analytical expected probability $\mu_{ij}^{d}$ that two nationalities $i$ and $j$ co-reside in a given district $d$ can be expressed as (Karimi et al.,, 2015):
$\mu_{ij}^{d}=N^{d}\left(N^{d}-1\right)P_{i}P_{j},$ (4)
where $N^{d}$ is the total population in the district $d$, and $P_{i}=\sum_{d}\kappa_{i}^{d}/N_{\text{City}}$ represents the proportion of residents of nationality $i$ relative to the total population of Vienna, $N_{\text{City}}=\sum_{d}N^{d}$.
To evaluate the significance of co-residence patterns, we calculate a standardised z-score for each nationality pair $(i,j)$ across districts, which indicates how much the observed co-residence deviates from the expected value under the null model. For a pair $(i,j)$ in district $d$, the z-score is defined as:
$z_{ij}^{d}=\frac{w_{ij}^{d}-\mu_{ij}^{d}}{\sigma_{ij}^{d}},$ (5)
where $\sigma_{ij}^{d}$ is the standard deviation of the null distribution for link weights. Using the multinomial theorem multiple times, one can compute the standard deviation as (Karimi et al.,, 2015):
$\left(\sigma_{ij}^{d}\right)^{2}=N^{d}\left(N^{d}-1\right)P_{i}P_{j}\left(\left(6-4N^{d}\right)P_{i}P_{j}+\left(N^{d}-2\right)\left(P_{i}+P_{j}\right)+1\right).$ (6)
The z-score standardises observed co-residence values, allowing comparisons between pairs regardless of population sizes or district densities. Finally, to capture the overall tendency of nationalities $i$ and $j$ to co-reside across Vienna, we sum z-scores over all districts to compute a cumulative z-score for each country pair:
$z_{ij}=\sum_{d}z_{ij}^{d}.$ (7)
This cumulative z-score allows us to assess the relative living tendency of nationality pairs in a way that accounts for variations across districts. Higher cumulative z-scores for a pair $(i,j)$ indicate a statistically significant co-residence pattern beyond random expectation, signifying a meaningful residence effect between the two nationalities. Very low z-scores would indicate a pattern of avoidance.
Clustering countries sharing similar residence interests.
To identify patterns of habitant clustering within Vienna, we examine how nationalities tend to co-reside across districts, revealing clusters of countries that share similar residential interests. By clustering these connections, we aim to uncover large-scale structures of co-residence patterns among national communities within the city.
To find these clusters, we employ a community detection approach that uses the strength of co-residence connections between nationalities across Vienna’s districts as a basis for grouping. The network of co-residence is constructed with each nationality pair $(i,j)$ linked by their co-residence weights $z_{ij}$, which reflect the frequency of their spatial overlap across districts (as detailed in the previous subsection and shown in Fig. 2b). This network provides the structural foundation to reveal groups of countries with significant shared residence patterns. Our approach to clustering relies on the network community-detection method known as the map equation, which leverages random walks to detect communities based on their flow within the network (Rosvall and Bergstrom,, 2008; Rosvall et al.,, 2009). In this context, the map equation algorithm simulates a sequence of “visits” to nationalities, where movement between nationalities is determined by the strength of their co-residence link. This method clusters nationalities based on how often a random walker would transition between them, prioritising strongly interconnected groups that are frequently co-resided within the same districts. Nationalities with stronger internal connections are likely to retain the walker for extended periods, which results in distinct, identifiable clusters. Here, we use the map equation’s associated search algorithm Infomap (Edler et al.,, 2023) to identify the groups of countries.
Notably, this method captures not only direct pairwise connections but also implicit relationships where two nationalities might share similar residence preferences due to common links with other nationalities, even if they do not directly co-reside frequently. The resulting clusters thus provide a simplified yet meaningful representation of the residential dynamics among national communities, offering a macroscopic perspective on their integration and spatial organisation within Vienna.
Influence of district diversity and national homophily on residential clustering.
Here, we leverage our dataset of migrants and locals in Vienna to investigate the inherent diversity within the city’s districts, as well as the degree of country-based homophily among residents. To achieve this, we apply two distinct statistical indices: the Simpson index and the Dissimilarity index. The Simpson index is used to quantify the diversity within each district, providing insight into the overall heterogeneity of the resident population in terms of nationality. A higher Simpson index reflects a more diverse population within a district, indicating a greater variety of nationalities living in close proximity. On the other hand, the Dissimilarity index is employed to measure homophily, or the tendency of individuals from the same nationality to cluster together within certain districts. This index helps us assess how much the population of each nationality tends to concentrate in specific areas rather than being evenly distributed across the city (Simpson,, 2007; Harris,, 2017). By analysing both the Simpson index for district diversity (see Fig. 4a and Fig. 4b) and the Dissimilarity index for homophily (see Fig. 4c and Fig. 4d), we gain valuable insights into the underlying reasons why certain nationalities in Vienna segregate into the two clusters identified in our analysis.
The Simpson index, which can be calculated for each district (local) and the entire city, is a measure of concentration. It expresses the probability that two randomly selected individuals from the district/city have different nationalities. Calculated for one district d of the city, the Simpson index, $S^{d}$, becomes (Duncan and Duncan,, 1955; White,, 1986):
$S^{d}=1-\sum_{i}\left(P_{i}^{d}\right)^{2},$ (8)
where $P_{i}^{d}=N_{i}^{d}/N^{d}$ is the fraction of inhabitants of nationality i residing in district d. It is a measure between 0 and a maximum value, which depends on the different number of nationalities, $C^{d}$, living in the district. The higher $S^{d}$, the more nationality-diverse the district is. A district is fully diverse ($S^{d}_{\text{max}}=1-1/C^{d}$) when the population in the district $d$ is equally distributed among all existing nationalities. In contrast, a district has no diversity ($S^{d}_{\text{min}}=0$), when it is inhabited exclusively by a single nationality. The Simpson index exhibits a close mathematical relationship with entropy, which serves as a quantitative measure of disorder (see Supplementary Note LABEL:sec:entropy). This relationship is often utilised within segregation studies to assess the unequal distribution of national groups across neighbourhoods or between global and local scales (Finney and Simpson,, 2009).
Furthermore, the city-wide Simpson index can be derived analogously to Eq. (8) by using the full population distribution in Vienna. The value for the city of Vienna is $S_{\text{City}}$ = 0.541. This number serves as a reference to compare with the average local Simpson index, which is the population-weighted average of the local Simpson index values in all districts. It is computed as (Zuccotti et al.,, 2023):
$S=\sum_{d}\left(\frac{N^{d}}{N_{\text{City}}}S^{d}\right).$ (9)
When the average local Simpson index is substantially lower than the city-wide Simpson index, it indicates a high degree of segregation, with districts tending to be highly homogeneous, often dominated by a single nationality. In Vienna, the average local Simpson index is $S$ = 0.535. This implies that there is approximately a 54% chance that a randomly selected pair of individuals from the same district will belong to different nationalities. This value closely aligns with the probability observed when sampling a pair randomly from the entire city, suggesting that overall segregation levels in Vienna are relatively low. These results are presented in Fig. 4a, where districts are ranked in descending order of diversity. The Simpson index is highest for district 10 and lowest for district 13 (see Supplementary Table LABEL:tab:tabs3 for detailed values). Additionally, Fig. 4b highlights the difference between each district’s diversity and the city-wide Simpson index, providing a clear comparison of which districts experience greater or lesser diversity relative to the city as a whole. We use this metric for calculating the Diversity-Population fraction correlation (Fig. 3b).
Furthermore, a detailed examination of Fig. 4b reveals a strong correlation with the percentage of migrants (see Fig. 1a) with a value corresponding to r[migrants, Diversity]=0.99 $\pm$ 0.01 (see the Supplementary Note LABEL:sec:diversity). It might initially seem that this pattern arises inherently from the definition of the Simpson index. However, if you consider a hypothetical scenario where all migrants in a specific district belong to a single nationality, the percentage of migrants would be high, yet the diversity would remain low. Therefore, this finding confirms that migrants are distributed relatively homogeneously across districts, residing in the more diverse ones and with no evidence of a specific migrant group forming a concentrated enclave in any particular district (see Fig. 1b and Supplementary Fig. LABEL:fig:figs4).
To define national homophily, the tendency to reside among others of the same nationality i, we calculate the Dissimilarity index $D_{i}$, which indicates the percentage of members of that group who would need to relocate so that the distributions at the district-level match the distribution at the city-level (Zuccotti et al.,, 2023):
$D_{i}=\frac{\sum_{d}\frac{N^{d}}{N_{\text{City}}}\cdot\left|P_{i}^{d}-P_{i}\right|}{2P_{i}\left(1-P_{i}\right)}.$ (10)
This index lies between 0 and 1. A nationality is fully segregated ($D_{i}=1$) when all the inhabitants of that origin reside in the same district. Conversely, a nationality is completely unsegregated ($D_{i}=0$) when its population is evenly distributed across all districts. The results are presented in Fig. 4c, where nationalities are ranked in descending order of homophily, and also illustrated in the world homophily map of Fig. 4d. The Dissimilarity index is highest for Turkish nationals and lowest for Hungarian (see Supplementary Table LABEL:tab:tabs4 for detailed values). There are notable differences among populations of various nationalities residing in Vienna, reflecting distinct residential patterns that provide insight into the dynamics of micro-level homophily. These differences underscore how individual preferences, rooted in shared language, traditions, and social norms, drive localised patterns of residence. Over time, these micro-level tendencies can scale up to shape the macro-level segregation seen in the city. For instance, even within broader cultural clusters, the presence of closely connected communities of a single nationality may amplify segregation effects, further dividing populations into distinct spatial and cultural domains. Understanding this interplay is essential, as it highlights the multi-layered nature of clustering, where both individual and group-level behaviours interact to produce the observed residential patterns.
We have also extensively examined several other indices using different statistical techniques in Supplementary Note LABEL:sec:othermeasures.
Figure 4: District diversity and national homophily in Vienna. a District diversity. Districts are ranked from highest to lowest based on the estimated diversity of Vienna districts. b Vienna diversity map. c National homophily. Nationalities are ranked from highest to lowest based on the estimated homophily of countries. d World homophily map.
Measuring correlations of the clustering key aspects.
The Pearson correlation coefficient, denoted as r, measures the strength and direction of a linear relationship between two continuous variables. It ranges from -1 to +1, where $r=+1$ indicates a perfect positive linear relationship, $r=-1$ denotes a perfect negative linear relationship, and $r=0$ suggests no linear association between the variables (Pearson,, 1895). Interpretation of r typically follows Cohen’s guidelines, which suggest that $|r|<0.3$ represents a weak relationship, $0.50.3\leq|r|<$ 0.5 indicates a moderate relationship, and $|r|\geq 0.5$ suggests a strong relationship (Cohen,, 2013). However, these cutoffs are context-dependent, and scientific judgment should consider the practical significance of the relationship in the specific field of study.
To compute r, we first assume two paired variables, X and Y, with n paired observations $(X_{i},Y_{i})$ for $i=1,2,\dots,n$. The formula for r is given by:
$r=\frac{\sum_{i=1}^{n}(X_{i}-\bar{X})(Y_{i}-\bar{Y})}{\sqrt{\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}}\sqrt{\sum_{i=1}^{n}(Y_{i}-\bar{Y})^{2}}},$ (11)
where $\bar{X}$ and $\bar{Y}$ are the sample means of X and Y, respectively. This equation standardises the covariation of X and Y by the product of their standard deviations, yielding a dimensionless value that facilitates comparison across different datasets and units (Fisher,, 1970).
To quantify the uncertainty around r, we calculate its standard error, $\sigma_{r}$, given by:
$\sigma_{r}=\sqrt{\frac{1-r^{2}}{n-2}},$ (12)
where n is the sample size. The standard error measures the extent to which the observed r might vary due to sampling variability, providing an estimate of its precision.
Pearson correlations shown in Fig. 3 were calculated for both subpanels a and b using the population fraction of each nationality relative to the total district population (see Supplementary Fig. LABEL:fig:figs4). In subpanel a, this population fraction was correlated with the average net income per district, derived from the additional dataset on income in Vienna’s districts (City of Vienna, 2023a, ) (see Supplementary Fig. LABEL:fig:figs2a). In subpanel b, the correlations were computed using the difference between each district’s diversity and the city-wide diversity (see Fig. 4b). Detailed Pearson correlation results for income and diversity are provided in Supplementary Table LABEL:tab:tabs4.
Data availability
The raw and processed data are not available due to privacy laws. The Federal Ministry of the Interior of Austria safeguarded the dataset and made it accessible to our research institution under strict data protection regulations. Researchers must reach individual agreements with the Federal Ministry of the Interior of Austria to access this data.
In GitHub repository (Parera,, 2024), we provide the aggregate statistics on the proportion of citizens of each nationality living in each district. Together with the publicly available complementary datasets, these data are sufficient to reproduce all the analyses performed in this study.
Code availability
The code developed to perform all analyses is available in GitHub repository (Parera,, 2024).
Acknowledgments
M.S. was partly funded by MCIN/AEI/ 10.13039/501100011033 and by “ESF Investing in your future”, grant number PRE2020-093266; by MCIN/AEI/ 10.13039/501100011033, by “ERDF A way of making Europe”, grant number PID2022-140757NB-I00, and the support of Generalitat de Catalunya, grant number 2021SGR00856. O.A. and R.P.C. were funded by the Austrian Federal Ministry for Climate Action, Environment, Energy, Mobility, Innovation and Technology (2021-0.664.668) and the Austrian Federal Ministry of the Interior (2022-0.392.231). F.K., S.M.G., and A.M.J. were partly funded by the EU Horizon Europe project MAMMOth (Grant Agreement 101070285). S.M.G. was supported by the Austrian research agency (FFG) under project No. 873927 ESSENCSE.
Author contributions statement
M.S. performed the analysis and developed models and simulations. M.S., S.M.G., A.M.J., and F.K. defined the problem and designed the solution and algorithms. O.A. and R.P.C. collected and processed the data. All authors designed research, analysed results, discussed results, and wrote and reviewed the manuscript.
Competing interests
The authors declare no competing interests.
Ethical approval
This article does not contain any studies with human participants performed by any of the authors.
Informed consent
This article does not contain any studies with human participants performed by any of the authors.